

********************************************************************
UARS WEEKLY STATUS REPORT - 22 FEB 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
SUSIM performed its daily mid-resolution and MgII solar scans
during periods of predicted UARS telemetry.
Measurements of the solar FUV irradiance were obtained for use in
comparing with TIMED/SEE and its calibration rocket by SUSIM at
mid (1.1 nm) and high (0.15 nm) resolutions on Feb. 8 and 11.
We ammend and correct last weeks report in the following. Most
backlogged level 0 data from December and January was obtained
directly from the CDHF. The deliveries of these same data from the
PACOR-A system are proceeding with recent deliveries of December
data. We are processing the data including those needed for the
MgII index as soon as they can be obtained irrespective of their
source.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the UARS ECLIPSE System
Meeting on 21 February 2002.
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 February 2002 (Orbit 56943, GMT Day 040) through
15 February 2002 (Orbit 57047, GMT Day 046).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 040/0005 56943 Power ON, ZEPS HEPS to day only
The beta angle ranged between the angles of 21.51 degrees to 39.49
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 08 March 2002 (DOY 067).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 24.80 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.31 to 0.63 degrees C. Additional
battery performance data for each battery on 09 February 2002 and 15
February 2002 are provided in Appendix D.
The clock error ranged from -13.0 to 12.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 8
PEM = 1
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 1
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/002 - 2287. 498845
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
RTS and ephemeris loads were sent to UARS from Eclipse as scheduled
this week. The RTS load was sent on 12 February 2002 on the 043/1755
and the Ephemeris load on the 045/1856 event. Both of the loads were
verified by telemetry and executed on board successfully. A
microprocessor load is scheduled to be sent to SUSIM next week.
MEETINGS
The FOT attended a UARS CECIL PROC Review on Monday, February 11,
2002.
The FOT attended a UARS Eclipse Working Meeting on Thursday, February
14, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 08 March 2002 (DOY 067). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 February 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
56943 040/0005 020-049 PEM Power ON, ZEPS HEPS to day only
56944 040/0225 02-050 OBC Clk Adj fr -20.6 to 0.00 msec/day
56986 042/2122 02-051 OBC Clk Adj fr 0.00 to 20.6 msec/day
56999 043/1755 02-052 OBC RTS 24 LOAD
57006 044/0410 02-052 OBC Clk Adj fr 20.6 to 0.00 msec/day
57012 044/1349 02-054 OBC Clk Adj fr 0.0 to -20.6 msec/day
57024 045/0919 02-055 OBC Clk Adj fr -20.6 to 0.00 msec/day
57027 045/1535 02-056 OBC Clk Adj fr 0.00 to 20.6 msec/day
57030 045/1856 02-057 OBC Ephem loaded through Eclipse system
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 February 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_____________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 February 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
56944 040/0236 Clk Adj fr -20.6 to 0.00 msec/day
56986 042/2124 Clk Adj fr 0.00 to 20.6 msec/day
57006 044/0411 Clk Adj fr 20.6 to 0.00 msec/day
57012 044/1352 Clk Adj fr 0.0 to -20.6 msec/day
57024 045/0920 Clk Adj fr -20.6 to 0.00 msec/day
57027 045/1539 Clk Adj fr 0.00 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 February 2002
09 February 2002 - GMT Day 040 Beta = 21.51 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -44.8/-67.2 -3.16/-4.08 +0.0/ -0.0 10.72 0.000 0.00
2 +22.4/-50.4 +6.47/+5.80 +29.6/-14.0 24.80 1.030 21.79
3 +22.4/-33.6 +5.47/+4.48 +29.2/-14.4 24.80 1.030 21.67
15 February 2002 - GMT Day 046 Beta = 39.49 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 -44.8/-56.0 -3.16/-4.38 +0.0/ -0.0 10.72 0.000 0.00
2 +11.2/-50.4 +5.14/+4.48 +27.2/-14.4 26.08 1.040 20.27
3 +22.4/-22.4 +4.48/+3.82 +26.4/-14.4 26.08 1.030 20.04
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 February 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-005 FOT 57016 044/2037 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-005 FOT see list Feb 02 14170 Generic Late Acquisition